OpenStudy (anonymous):

2(4x + 3) - 9 = 45 distribute then solve for x

OpenStudy (tennistar):

go t the qustion underneath you and see how ti was solved

mathslover (mathslover):

\[\huge{a(b+c) =ab+ac}\] \[\huge{2(4x+3)=2(4x)+2(3)=8x+6}\]

OpenStudy (anonymous):

8x+6-9=45 8x=48 x=6

mathslover (mathslover):

8x+6 - 9 = 8x-3

OpenStudy (anonymous):

First we would do distribution. That would give us 8x+6-9=45. Now we combine the numbers on the left and get 8x-3=45. Once we have this, we add both sides by 3 and get 8x=48. Finally we divide by 8 and get x=6.

mathslover (mathslover):

now just equate this to 45 8x-3=45 8x=48 x=\(\frac{48}{8}\) = 6

mathslover (mathslover):

Note what i did here is very easy : break down the equation to make it easier : this is the given equation right ? 2(4x + 3) - 9 = 45 now break this into 2(4x+3) and subtract 9 and equate it as 45

OpenStudy (anonymous):

2(4x +3) -9 =45 8x-3=45 8x =48 (adding 3 on both sides) x=6 (dividing by 8 on both sides)